<e> A vending treasury is purchased for RM 2,800 in cash with RM 700 down payment and fourteen monthly payments. If the interest calculated is 10% per annum, using the fixed rate equation calculation find:


1. Total interest charged


2.Secondly. monthly payment



1
Expert's answer
2022-02-22T00:58:36-0500

Amount need to borrow = 2800-700 = 2100

"Monthly Rate = \dfrac{10\%}{12} = 0.8333\%"

"Monthly Payment = \dfrac{Amount\ need\ to\ borrow}{\dfrac{1-(1+r)^{-n}}r}=\dfrac{2100}{\dfrac{1-(1+0.8333\%)^{-14}}{0.8333}}=159.54\$"

Total interest charged = Amount to be paid in total - Amount Borrowed

"=14\times159.54-2100"

=$133.60